OpenStudy (anonymous):

Use the substitution meth x + y = 3 y = 9 A. (6, 9) B. (-6, 9) C. (-12, 9) D. (12, 9)

Parth (parthkohli):

B. (-6,9)

Parth (parthkohli):

Can I explain?

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

B is the correct one. plug in y=9 in x+y=3 and u get x= -6

Parth (parthkohli):

We know that y = 9(2nd line) And the first line says the x + y = 3 So, x + 9 = 3 x = 3-9 x = -6